OpenStudy (anonymous):

Solve the following system of equations : 2y=8-x y-2x=-1 Solution : ( ) , ( )

OpenStudy (anonymous):

i need help on this onee

OpenStudy (anonymous):

Rearrange the second equation to yield y=2x - 1 Substitute this into the first equation to yield 2(2x-1)=8-x you can then rearrange this to find an expression for x; x=2 Finally, substitute this into the second equation to find y; y=3

OpenStudy (anonymous):

Thanks so much alde.e

OpenStudy (anonymous):

2y=8-x y-2x=-1 x=0 x=8 x=0 x1/2 y=4 y=0 y=-1 y=0